OpenStudy (anonymous):

How do I solve for x if I have x^(-5/2)=(8/3)?

OpenStudy (kinggeorge):

\[{x^{-{5 \over 2}}} = {{1 \over x^{5 \over 2}}} = {1 \over \sqrt{x^5}} = {8 \over 3}\]so\[\sqrt{x^5} = {3 \over 8} \longrightarrow {x^5}= {9 \over 64}\]so\[x = \sqrt[5]{9 \over 64}\]

OpenStudy (anonymous):

Thanks!